Session: Mar 2007 (session 1)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (G.Maccaferri at Wed Mar 21 09:03:54 2007) Minor failures : RFI

Westerbork (Tony Foley at Wed Mar 21 09:09:39 2007) Success :
Single dish was used. Please change the coordinates in the vexfile accordingly. Looks fine, but some of our automatic scripts to check progress were unavailable (server problems)

Torun (Kaz Borkowski at Wed Mar 21 09:45:16 2007) Minor failures : RECE
Only one channel recorded (see notes for F07M1 and N07M1). In addition, 7 minutes lost starting with the second scan due to unwrapping of cables in azimuthal telescope position.

Jodrell_Bank (Alastair Gunn at Wed Mar 21 11:51:10 2007) Success :
The MkII telescope was used (Jb2). No known problems.

Cambridge (Alastair Gunn at Wed Mar 21 11:51:25 2007) Success :
No known problems.

Onsala (Michael Lindqvist at Wed Mar 21 13:07:49 2007) Success :
No known problems.

Effelsberg (A. Polatidis at Fri Mar 23 09:44:11 2007) Success :

Noto (P. Cassaro at Sun Mar 25 23:29:34 2007) Success :